# Towards AI-Assisted Synthesis of Verified Dafny Methods

### v1

@@ Abstract.
Large stochastic language models show great promise in many domains, including programming. A promise is easy to make but hard to keep, and language models often fail to keep their promises when applied to programming, generating erroneous code. One promising avenue to keep models honest is to have them generate code in a language that supports formal verification: if and when that is adopted, the model would provide proof along with the code, and that proof would be automatically verified. Unfortunately, existing large language models show a severe lack of proficiency in verified programming languages.
In this paper we demonstrate how to improve two pretrained models’ proficiency in the Dafny verified programming language. Using 178 programming problems from the MBPP dataset, we prompt two contemporary models (GPT-4  and PaLM-2 ) to generate methods in Dafny. We use three different types of prompts: a direct contextless prompt, a second one that includes a signature of the method and test cases, and a third one that decomposes the problem into steps and includes dynamically chosen similar examples. Our results show that GPT-4 is better than PaLM-2 , but that, in both models, the third prompt greatly improves the success of the generation task with respect to the direct prompt. With the third prompt, GPT-4 was able to generate verified (and human-evaluated) Dafny methods in 58% of the cases, while the first prompt generated verified (and human-evaluated) methods in only 19% of the cases. Surprisingly, the second prompt had the worst performance, with only 10%.
One tangible contribution of our work is a collection of 153 MBPP problems that are implemented and formally verified in Dafny, 50 of which were written by us and 103 were automatically synthesized by GPT-4 . Additionally, our results demonstrate that the benefits of formal program verification (proof of correctness) are now within reach of large stochastic language models used to generate code. These results also demonstrate that program verification systems can likewise benefit from incorporating large language models, whether to synthesize code wholesale, to generate specifications, or to construct internal verification annotations such as loop invariants, that are hard for programmers and verification tools to find directly. (e.g. legal arguments, transport signaling, structural engineering, etc.) where solutions must be correct, and where that correctness needs to be verifiable by existing formal tools, or explained to (and understood by) designers and end-users.
@@ 1. Introduction
Formal verification is increasingly being adopted to support the development of high-quality, provably correct software. High-stakes domains, such as security-sensitive systems, cryptographic libraries, aerospace systems, and embedded software in medical devices, look to formal verification for correctness guarantees.
@@
For instance, CompCert C (CompCert, 2023) is a formally verified C compiler that won the ACM Software Systems award in 2021;  [^1] https://awards.acm.org/software-system in a comprehensive study (Yang et al., 2011), researchers found no bugs in the CompCert C compiler compared to GCC (GCC, 2023) and LLVM (LLVM, 2023) toolchains. This study motivated Airbus to adopt CompCert C to help ensure safety and enhance aircraft performance (França et al., 2011). The seL4 project (Klein et al., 2009), awarded the ACM Software System Award in 2022, resulted in a formally verified high-assurance, high-performance operating system microkernel employed to protect an autonomous helicopter against cyber-attacks (seL4 Project, 2023). In order to secure communication, both Chrome and Android use formally verified cryptographic code (Erbsen et al., 2020). Similarly, Mozilla incorporated its verified cryptographic library for Firefox performance improvement (Jacobs and 2020, 2023).
@@
Formal code verification consists of two parts that go hand-in-hand: formal specification of software properties, and automated, or semi-automated, verification of those properties. Over the past 50 years, verification has had a couple of major breakthroughs, first with the development of interactive theorem provers in the 1960s (Nederpelt et al., 1994), and then, at the turn of the millennium, with the development of Satisfiability Modulo Theory (SMT) solvers (Barrett et al., 2021). Unfortunately writing program properties and proofs is still a creative, manual process that requires significant effort, experience, and expertise. Formal specification languages are closer to mathematics than to regular programming languages, and necessarily incorporate subtle concepts based on classical or constructive logics. For example, writing the proofs for the seL4 microkernel was an eleven person-year effort of a group of experts (Murray et al., 2013). Another example: the verification code of the CompCert C compiler is more than three times the size of the original compiler code itself; it took three person-years to write the complete verification code, which is approximately two lines of code in a day (Leroy, 2009). The fact that these systems have received important awards is a testament to how exceptional these efforts are.
@@
In the universe of languages and systems for formal code verification, Dafny (Microsoft, 2023a) stands out as having made a significant effort towards usability by programmers, rather than mathematicians. Dafny is a strongly typed imperative programming language with functional and object-oriented features that supports code verification via Hoare Logic (Hoare, 1969) assertions, preconditions, postconditions, and invariants (aka design by contract). Although algorithmic code in Dafny is similar to many other programming languages, writing the formal specifications and auxiliary verification assertions is still difficult (Faria and Abreu, 2023; Noble et al., 2022).
@@
Over the past year, stochastic Large Language Models (LLMs) have been revolutionizing both the software industry and research in software engineering. LLM-assisted tools such as GitHub-Copilot (Copilot, 2023) and Amazon CodeWhisper (CodeWhisperer, 2023) have been accelerating several tasks of software development such as code generation, transformation, summarization, documentation, code review, program repair and synthesis. Although their performance is impressive for straightforward tasks in the most popular programming languages, given more complex problems or less familiar languages (such as Dafny), LLMs tend towards producing ”hallucinations” [^2] Respectable software engineers would call hallucinations ”faults”, ”errors”, or ”bugs”. (McKenna et al., 2023; Galitsky, 2023; Salvagno and Taccone, 2023; Azaria and Mitchell, 2023) — inventing control structures or APIs that do not exist in the target programming language, borrowing features from other languages, often generating code that fails to type-check or even fails to parse.
@@
  • We conduct the first empirical study of LLMs synthesizing verifiable Dafny methods.
  • We demonstrate that a prompt designed under the principles of Chain of Thought (CoT) with semantically similar few shot examples explaining how to decompose a problem step by step, can synthesize verified and correct Dafny methods with meaningful specifications for 58% of problems in our test dataset.
  • We release the dataset MBPP-DFY-153, a collection of 153 programming problems with specifications, solutions and tests in/for Dafny, based on the MBPP (Mostly Basic Python Programming) dataset curated by Google (Austin et al., 2021).
  • We demonstrate that program verification techniques can help ensure code generated by LLMs is correct.
Our conclusions lead to a number of actionable insights. Most obviously, our first contribution means that programmers developing Dafny systems could already consider using LLMs — indeed, for simple cases, an LLM may be able to generate both the specificiations of a Dafny method, and a verifiably correct method implementation that meets those specifications. For researchers in LLMs or verification, our second contribution, the dataset of 153 formally verified Dafny examples, should enable further development of prompts and models to improve the performance of LLMs.
@@ 2.1. Dafny: A Verification-Aware Programming Language
Dafny (Microsoft, 2023a; Leino, 2020) is a verification-aware statically typed programming language which was first developed in Microsoft Research (MSR) (Microsoft, 2023b); it is currently being developed with the support of the Amazon Automated Reasoning research group (Amazon, 2023). Dafny includes various features of several programming paradigms such as imperative, functional, and object oriented programming. Its toolchain includes translators for generating executable code in different target languages such as C#, Java, JavaScript, Go and Python (Dafny, 2023). But the distinguishing feature of Dafny is that it supports code verification via design by contract. For that, it follows the framework of Floyd-Hoare-style  (Hoare, 1969) program verification with preconditions, postconditions, loop invariants and other high-level formal proof synthesis features such as pure functions, predicates, lemmas, and automated proof by inductions.
To develop a verified program, developers write Dafny code along with the specifications (pre-and post-conditions), and then add loop invariants and assertions that help the Dafny verifier prove the correctness of their code. While coding in VS Code, the Dafny static program verifier checks the functional correctness based on developers’ defined specifications and annotations. Correctness means the absence of any runtime errors with respect to the formal pre- and post-conditions, which means that the code does what the developers specify it to do. In order to confirm that the developer
1 : method FindSmallest(s: array<int>) returns (min: int)
2 :     requires s.Length > 0  // preconditions
3 :     // postconditions
4 :     ensures forall i :: 0 <= i < s.Length ==> min <= s[i]
5 :     ensures exists i :: 0 <= i < s.Length && min == s[i]
6 : {
7 :     min := s[0];
8 :     for i := 1 to s.Length
9 :         invariant 0 <= i <= s.Length  //loop-invariants
10:         invariant forall k :: 0 <= k < i ==> min <= s[k]
11:         invariant exists k :: 0 <= k < i && min == s[k]
12:     {
13:         if s[i] < min { min := s[i]; }
14:     }
15: }
16: //Main program, with a test case (checked statically!)
17: method Main() {
18:     var a1:= new int[] [1,2,3];
19:     var out1:= FindSmallest(a1);
20:     assert out1 == 1;
21: }
Figure 1. An example Dafny method; FindSmallest
specifications holds, the Dafny program verifier first transforms the code into an intermediate verification representation (IVR) known as Boogie IVR (Le Goues et al., 2011) language that expresses the verification conditions into predicate calculus (Leino, 2017). Next, an SMT solver tries to prove the verification conditions. In this phase, the Dafny verifier is backed by an advanced SMT solver known as the Z3 theorem prover (de Moura and Bjørner, 2008). The validity of these verification conditions implies the correctness of the code under consideration (Leino, 2017). However, sometimes the Z3 solver cannot automatically reach the required proof, even though such proof exists. In such cases, the developer intervention is required to give more context in the specifications by writing auxiliary verification conditions in the form of assertions, functions, predicates, and lemmas.
@@ 2.2. Motivating Example
To illustrate how a Dafny program is statically verified, Figure 1 shows a simple procedure to find, and return, the smallest number in a given input array. The method FindSmallest is formally specified not just by the types of input and output parameters, but also with preconditions (with requires clause) and postconditions (with ensures clause). The postconditions, in particular, are the properties that we want to ensure are true for this method – i.e. what it means for this method to be “correct.” The Z3 SMT solver is usually not powerful enough to infer the postconditions from functional code alone; as such, developers must give additional hints, or verifiable conditions, along with the code in order to help the solver prove correctness of the postconditions.
In this particular case, the code contains a loop that modifies state, so we can provide loop invariants (with invariant clause) that must hold in every iteration of the loop; those invariants
are then checked by the SMT solver and, if they hold, that helps proving the postconditions of the method. The Main method represents the entry point: it invokes the method FindSmallest for some test inputs and checks (with assert clause) the outputs. Like other specification clauses, the assert statements are also statically checked. In this example the verifier tries to prove the assertion based on the postconditions of the FindSmallest method. All specification clauses such as requires, ensures, invariants, and assert are used statically – these clauses are not compiled into the executable and, as such, do not cause any runtime overhead.
Writing the functional code in Dafny is similar to writing code in other programming languages. However, researchers observed that writing the formal specifications and auxiliary verification code in Dafny is still difficult (Faria and Abreu, 2023). For instance, in Figure 1, the FindSmallest method requires 6 lines of formal specifications to verify a simple 5 line of algorithmic code. This example substantiates that the difficulty of writing formal specifications exponentially increases with the difficulty of the verifying program (Noble et al., 2022).
@@ 2.3. Large Language Models (LLMs)
Large Language Models (LLMs) are a class of deep neural networks pre-trained in publicly available data to perform a variety of downstream tasks in natural language processing (NLP). In the past year, LLMs have significantly impacted a wide range of application domains including code related tasks such as code generation, comprehension, summarization, and transformation.
The recent improvement of transformer based models (Radford et al., 2018) such as GPT (Generative Pre-trained Transformer) advanced the usages of LLMs in many coding related tasks compared to other pre-trained models (Chen et al., 2023; OpenAI, 2023a). For instance, CodeX  (Mark Chen et al., 2021) is a descendant of the GPT-3 based model trained on massive publicly available code repositories in GitHub, and is capable of performing coding related tasks in various programming languages. Codex powers GitHub Copilot (Copilot, 2023), an AI-pair programmer designed to assist software developers. OpenAI (OpenAI, 2023b) has now discontinued its Codex APIs, and suggested researchers adopt the successors of GPT-3 (GPT-3.5 and GPT-4). Recently Google released PaLM-2 (Rohan Anil et al., 2023) a new successor of their earlier PaLM (Aakanksha Chowdhery et al., 2022) model which is very capable of code generation and reasoning tasks.
Although many research studies have been published on the use of LLMs in many coding-related tasks, very few of them explore these models’ capability for synthesis of formally verified programs (First et al., 2023). One of the obstacles is that languages for formal code verification are not very popular. Hence, the training data available for LLMs to learn these languages, and their underlying concepts, pales when compared to the amount of training data available for popular languages such as Python, Java, C++, etc. Nevertheless, LLMs have been found to be capable of working as generic reasoning engines, including having the capability to solve complex problems and prove theorems (Li et al., 2023; Wu et al., 2022) that motivates us exploring how LLMs with the combination of theirs reasoning, theorem proving, and coding abilities can contribute to synthesizing verified Dafny code, along with formal specifications and validation conditions.
@@ 2.4. Prompt Engineering
Prompt engineering is a relatively new term and concept that emerged in the context of LLMs. It denotes the practice of iteratively crafting prompts for improving the generative power and fitness of purpose of an LLM. First coined by Gwern Branwen in 2020, [^3] https://gwern.net/gpt-3 prompt engineering has become a thriving subfield of research in LLMs (Bach et al., 2022; Dang et al., 2022; Deckers et al., 2023; Hou et al., 2022; Jiang et al., 2022b; Liu and Chilton, 2022).
LLMs are nominally intuitive to interact with: we give them a prompt, and they reply with a related text or image. The resulting output of the model relates to our prompt in plausible form. However, when using LLMs for specific technical goals, the simplistic approach of asking for something informally usually ends up in disappointing results.
Prompt engineering involves a systematic approach of creating prompts that effectively convey the desired downstream task to the pre-trained LLMs (Reynolds and McDonell, 2021; Beurer-Kellner et al., 2023). The crafting of prompts plays a significant role in the performance of LLMs retrieved contents. The prompt template should be structured in a manner that aligns with both the targeted task and the selected LLMs. Prompting techniques can be employed to train the LLMs explicitly through few-shot learning, incorporating a limited number of examples, such as method signatures and test cases. In the context of formally verified programs synthesis, the prompts need to be tailored for the targeted programming language (in our case Dafny) with unambiguous task descriptions (Bach et al., 2022; Dang et al., 2022).

### v2

@@ Abstract.
Large language models show great promise in many domains, including programming. A promise is easy to make but hard to keep, and language models often fail to keep their promises, generating erroneous code. A promising avenue to keep models honest is to incorporate formal verification: generating programs’ specifications as well as code, so that the code can be proved correct with respect to the specifications. Unfortunately, existing large language models show a severe lack of proficiency in verified programming.
In this paper, we demonstrate how to improve two pretrained models’ proficiency in the Dafny verification-aware language. Using 178 problems from the MBPP dataset, we prompt two contemporary models (GPT-4 and PaLM-2) to synthesize Dafny methods. We use three different types of prompts: a direct Contextless prompt; a Signature prompt that includes a method signature and test cases, and a Chain of Thought (CoT) prompt that decomposes the problem into steps and includes retrieval augmentation generated example problems and solutions. Our results show that GPT-4 performs better than PaLM-2 on these tasks, and that both models perform best with the retrieval augmentation generated CoT prompt. GPT-4 was able to generate verified, human-evaluated, Dafny methods for 58% of the problems, however GPT-4 managed only 19% of the problems with the Contextless prompt, and even fewer (10%) for the Signature prompt. We are thus able to contribute 153 verified Dafny solutions to MBPP problems, 50 that we wrote manually and 103 synthesized by GPT-4.
Our results demonstrate that the benefits of formal program verification are now within reach of code-generating large language models. Likewise, program verification systems can benefit from large language models, whether to synthesize code wholesale, to generate specifications, or to act as a ”programmer’s verification apprentice”, to construct annotations such as loop invariants which are hard for programmers to write or verification tools to find. Finally, we expect that the approach we have pioneered here — generating candidate solutions that are subsequently formally checked for correctness — should transfer to other domains (e.g., legal arguments, transport signaling, structural engineering) where solutions must be correct, where that correctness must be demonstrated, explained and understood by designers and end-users.
@@ 1. Introduction
Formal verification is increasingly being adopted to support the development of high-quality, provably correct software. High-stakes domains, such as security-sensitive systems, cryptographic libraries, aerospace systems, and embedded software in medical devices, look to formal verification for correctness guarantees. For instance, CompCert C (CompCert, 2023) is a formally verified C compiler that won the ACM Software Systems award in 2021111https://awards.acm.org/software-system: in a comprehensive study (Yang et al., 2011), researchers found no bugs in the CompCert C compiler compared to the GCC (GCC, 2023) and LLVM (LLVM, 2023) tool-chain. This study motivated Airbus to adopt CompCert C to help ensure safety and enhance aircraft performance (França et al., 2011). The seL4 project (Klein et al., 2009), awarded the ACM Software System Award in 2022, resulted in a formally verified high-assurance, high-performance operating system microkernel employed to protect an autonomous helicopter against cyber-attacks (seL4 Project, 2023). In order to secure communication, both Chrome and Android use formally verified cryptographic code (Erbsen et al., 2020). Mozilla incorporated its verified cryptographic library for Firefox performance improvement (Jacobs and Beurdouche, 2022).
@@
Formal code verification consists of two parts that go hand-in-hand: formal specification of software properties, and automated, or semi-automated, verification of those properties. Over the past 50 years, verification has made a number of significant breakthroughs, first with the development of interactive theorem provers in the 1960s (Nederpelt et al., 1994), and then, at the turn of the millennium, with the development of Satisfiability Modulo Theory (SMT) solvers (Barrett et al., 2021). Unfortunately writing program properties and proofs is still a creative, manual process that requires significant effort, experience, and expertise. Formal specification languages are closer to mathematics than to regular programming languages, and necessarily incorporate subtle concepts based on classical or constructive logics. For example, writing the proofs for the seL4 microkernel was an eleven person-year effort of a group of experts (Murray et al., 2013). The additional code required to verify the CompCert C compiler is more than three times the size of the compiler itself; and took three person-years to write, approximately two lines of code per day (Leroy, 2009). That these systems have received important awards is a testament to the exceptional effort required.
@@
In the universe of languages and systems for formal code verification, Dafny (Microsoft, 2023a) stands out as having made a significant effort towards usability by programmers. Dafny is a strongly typed imperative programming language with functional and object-oriented features that supports code verification via Hoare Logic (Hoare, 1969) assertions, preconditions, postconditions, and invariants — often known as Design by Contract (Meyer, 1992). Although algorithmic code in Dafny is similar to many other programming languages, writing the formal specifications and auxiliary verification assertions is still difficult (Faria and Abreu, 2023; Noble et al., 2022).
@@
Over the past year, Large Language Models (LLMs) have been revolutionizing both the software industry and research in software engineering. LLM-assisted tools such as GitHub’s Copilot (Copilot, 2023) and Amazon CodeWhisper (CodeWhisperer, 2023) have been applied to many software development tasks, including as code generation, transformation, summarization, documentation, code review, program repair, and synthesis. Although their performance is impressive for straightforward tasks in popular programming languages, given more complex problems or less familiar languages (such as Dafny), LLMs tend to produce ”hallucinations” [^2] Respectable software engineers call hallucinations ”faults”, ”errors”, or ”bugs”. (McKenna et al., 2023; Galitsky, 2023; Salvagno and Taccone, 2023; Azaria and Mitchell, 2023) — inventing control structures or APIs that do not exist in the target programming language, borrowing features from other languages, and often generating code that fails to type-check or even fails to parse.
@@
[…]
Our conclusions lead to a number of actionable insights. Most obviously, our first contribution means that programmers developing Dafny systems could already consider using LLMs — indeed, for simple cases, an LLM may be able to generate both the specifications of a Dafny method, and a verifiably correct method implementation that meets those specifications. For researchers in LLMs or verification, our second contribution, the dataset of 153 formally verified Dafny examples, should enable further development of prompts and models to improve the performance of LLMs.
@@ 2.1. Dafny: A Verification-Aware Programming Language
Dafny (Microsoft, 2023a; Leino, 2023) is a verification-aware statically typed programming language which was first developed in Microsoft Research (MSR) (Microsoft, 2023b); it is currently being developed with the support of the Amazon Automated Reasoning research group (Amazon, 2023). Dafny includes features drawn from programming paradigms including imperative, functional, and object-oriented programming. Dafny’s tool-chain includes translators for generating executable code in several target languages including C#, Java, JavaScript, Go and Python (Dafny, 2023). The distinguishing feature of Dafny is that it supports code verification via design by contract. For that, it follows the framework of Floyd-Hoare-style  (Hoare, 1969) program verification with preconditions, postconditions, loop invariants and other high-level formal proof synthesis features such as pure functions, predicates, lemmas, and automated proof by induction.
To develop a verified program, developers write Dafny code along with the specifications (pre-and post-conditions), and then add loop invariants and assertions that help the Dafny verifier prove the correctness of their code. While coding in VS Code, the Dafny static program verifier checks the functional correctness based on developers’ defined specifications and annotations. Correctness means the absence of any runtime errors with respect to the formal pre- and postconditions, which means that the code does what the developers specify it to do. In order to confirm that methods’ specifications hold, the Dafny program verifier first transforms the code into an intermediate verification representation (Goues et al., 2011) that encodes the verification conditions in predicate calculus (Leino, 2017), and then invokes the Z3 SMT solver (de Moura and Bjørner, 2008) to prove the verification conditions. The validity of these verification conditions implies the correctness of a program’s code (Leino, 2017). Z3 cannot always prove the verification conditions, even where such a proof exists. In these situations, developers need to guide the solver, e.g. giving more context by writing auxiliary verification annotations such as assertions, functions, predicates, and lemmas.
@@ 2.2. Motivating Example
To illustrate how a Dafny program is statically verified, Figure 1 shows a simple procedure to find, and return, the smallest number in a given input array. The method FindSmallest is formally
Find the smallest element in an array 1 : method FindSmallest(s: array<int>) returns (min: int) 2 : requires s.Length > 0 // preconditions 3 : // postconditions 4 : ensures forall i :: 0 <= i < s.Length ==> min <= s[i] 5 : ensures exists i :: 0 <= i < s.Length && min == s[i] 6 : { 7 : min := s[0]; 8 : for i := 1 to s.Length 9 : invariant 0 <= i <= s.Length //loop-invariants 10: invariant forall k :: 0 <= k < i ==> min <= s[k] 11: invariant exists k :: 0 <= k < i && min == s[k] 12: { 13: if s[i] < min { min := s[i]; } 14: } 15: } 16: //Main program, with a test case (checked statically!) 17: method Main() { 18: var a1:= new int[] [1,2,3]; 19: var out1:= FindSmallest(a1); 20: assert out1 == 1; 21: }
Figure 1. An example Dafny method; FindSmallest
specified not just by the types of input and output parameters, but also with preconditions (with requires clause) and postconditions (with ensures clause).
The postconditions are the properties we want to ensure are true when this method returns: what it means for this method to be “correct.” The Z3 SMT solver is not always able to infer postconditions, especially across imperative code. Figure 1 contains a loop that modifies state: we must provide loop invariants that hold before and after every iteration of the loop.
The Main method is the entry point to the program; it invokes the method FindSmallest for some test inputs and checks (with assert) the outputs. Like other specifications in Dafny, assert statements are also statically checked and do not incur any runtime overhead. Writing Dafny code is similar to writing code in other imperative and object-oriented programming languages, however, researchers observed that writing formal specifications and auxiliary verification annotations in Dafny is still difficult (Faria and Abreu, 2023; Noble et al., 2022). We can see this in Figure 1, which requires 6 lines of formal annotations to verify the 5 lines of the FindSmallest method body.
@@ 2.3. Large Language Models (LLMs)
Large Language Models (LLMs) are a class of deep neural networks trained on publicly available data to perform a variety of downstream tasks in natural language processing (NLP) (Madaan et al., 2022). In the past year, LLMs have significantly impacted a wide range of application domains including code related tasks such as code generation, comprehension, summarization, and transformation (First et al., 2023; Li et al., 2023; Wu et al., 2022).
The recent improvement of transformer based models (Radford et al., 2018) such as GPT (Generative Pre-trained Transformer) advanced the use of LLMs in many coding related tasks compared to other pre-trained models (Chen et al., 2023; OpenAI, 2023a). For instance, CodeX  (Mark Chen et al., 2021) is a descendant of GPT-3 trained on massive publicly available code repositories in GitHub, and is capable of performing coding related tasks in various programming languages. Codex powers GitHub Copilot (Copilot, 2023), an AI “programmer’s assistant” designed to assist software developers. OpenAI (OpenAI, 2023b) has now discontinued its Codex APIs, and suggested researchers adopt the successors of GPT-3 (GPT-3.5 and GPT-4). Recently, Google released PaLM-2 (Rohan Anil et al., 2023), a successor to their earlier PaLM (Aakanksha Chowdhery et al., 2022) model which is capable of code generation and reasoning tasks.
Unfortunately, the applicability of LLMs to formally verified programming languages has not been studied extensively  (First et al., 2023). Languages that support formal verification are not popular, so there is a paucity of training data available. LLMs have, however, been found to be capable of working as generic reasoning engines solving complex problems and proving theorems (Li et al., 2023; Wu et al., 2022). These recent successes motivated us to explore how contemporary LLMs can contribute to formal software development in Dafny.
@@ 2.4. Prompt Engineering
Prompt engineering (Bach et al., 2022; Dang et al., 2022; Deckers et al., 2023; Hou et al., 2022; Jiang et al., 2022b; Liu and Chilton, 2022), coined by Gwern Branwen in 2020333https://gwern.net/gpt-3, describes the incremental and iterative process of crafting inputs (“prompts”) for LLMs. LLMs commonly support conversational interaction: we give them a prompt, and they reply with a related text or image, plausibly derived from the prompt. To apply LLMs to specific technical tasks, however, an informal, simplistic approach to prompting usually leads to disappointing results. Effective prompts need to be structured to align with both the targeted task, and the LLM that will receive the prompt.
Prompt engineering takes a systematic approach to creating prompts to convey tasks to pre-trained LLMs (Reynolds and McDonell, 2021; Beurer-Kellner et al., 2023). To synthesize formally verified programs, prompts need to be engineered so that task descriptions are as unambiguous as possible (Bach et al., 2022; Dang et al., 2022), and solution examples must suit the target programming language (in our case Dafny). Prompt engineering techniques can also train LLMs explicitly, e.g.  through few-shot learning with a limited number of example problems and solutions: in our domain, textual problem description, method signature, and test cases.